could anyone help me with this qustion please.
find delta y/delta x, give the following iformation.
a) y=1/x, x=3, delta x=0.01
b) y=squrut x, x=9, delta x=0.9

Answer by Alan3354(69443)

find delta y/delta x, give the following iformation.
a) y=1/x, x=3, delta x=0.01
at x = 3, y = 1/3
at x = 3.01, y = 1/3.01
delta y = 1/3.01 - 1/3 = (3-3.01)/3*3.01 = -0.01/9.03
dy/dx = (-0.01/9.03)/0.01
dy/dx = -1/9.03
dy/dx =~ -0.11074
---------------
b) y=sqrt x, x=9, delta x=0.9
At x = 9, y = 3
At x = 9.9, y = sqrt(9.9) = 3sqrt(1.1)
dy = 3-3sqrt(1.1)
dy/dx = (3-3sqrt(1.1))/0.9
dy/dx = (1-sqrt(1.1))/0.3
dydx =~ -0.162696